The length of a rectangle is 2yd longer than it's width. If the permitter is 32yd find it's area.

Answer by mananth(16949)

The length of a rectangle is 2yd longer than it's width. If the permitter is 32yd find it's area.
Let width be x
Length = x+2 yards
Perimeter = 2*(l+w)
= 2*(x+2+x)
=2*(2x+2)
=4x+4
4x+4 = 32
4x=32-4
4x=28
x=7 = width
length =9

mananth@hotmail.com